OpenStudy (anonymous):

I need help with algebra 1 honors!!!! Please!

OpenStudy (anonymous):

What's the question??

OpenStudy (lgbasallote):

would be nice if you would be specific on which part of algebra 1 you're having a hard time...@StealerGirly50

OpenStudy (anonymous):

1. [9.01] When the quadratic equation y = (x − 5)(2x + 3) is written in standard form, what is the value of the coefficient "b"? (4 points) 7 −7 2 −15 2. [9.01] Identify which equation is a quadratic equation. (4 points) y = −5x − 3 y = −4x(x + 10) + 34 y + x2 = (x − 4)(x + 3) y − 2 = (x − 2)(x2 + 1) 3. [9.01] Determine whether the graph of y = −3x2 − 4x + 10 opens up or down and whether it has a maximum or minimum point. (4 points) Down and Minimum Down and Maximum Up and Maximum Up and Minimum 4. [9.01] What is the vertex for the graph of y = −2x2 + 8x − 13? (4 points) (2, 11) (2, −5) (−2, −37) (−2, −21) 5. [9.01] The graph of which equation is shown below? (4 points) y = −x2 + 6x − 5 y = −x2 − 6x − 5 y = x2 + 6x − 5 y = x2 − 6x − 5 6. [9.02] What are the solutions of x2 − 3x − 18 = 0? (4 points) x = 3, x = −6 x = −3, x = 6 x = 3, x = 6 x = −3, x = −6 7. [9.02] What are the x-intercept(s) of the graph of y − x2 = −3x − 10? (4 points) (2, 0) and (−5, 0) (−2, 0) and (5, 0) (−2, 0) and (−5, 0) (2, 0) and (5, 0) 8. [9.02] What are the solutions of 6x2 − x − 2 = 0? (4 points) x = , x = − x = −, x = x = , x = −1 x = −, x = 2 9. [9.02] Solve 6x2 = 18x. (4 points) x = 0, x = −3 x = 0, x = 3 x = 6, x = −3 x = 6, x = 3 10. [9.01], [9.02] Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x2 − 25 = 0 and x2 = 2x + 15 Part 2: Describe what the solution(s) represent to the graph of each. (2 points) Part 3: How are the graphs alike? How are they different? (2 points) (7 points) 11. [9.01], [9.02] Create your own unique quadratic equation in the form y = ax2 + bx + c. Use complete sentences and show all work to determine the following: Does the graph open up or down? How do you know? (2 points) Explain whether the graph has a maximum or minimum point. (2 points) Find the vertex and x-intercepts of the graph. (3 points) (7 points)

OpenStudy (anonymous):

I am pretty sure on some of them. I want to double check.

OpenStudy (anonymous):

You dont have to answer the short answers just tell me how to work the questions out

OpenStudy (anonymous):

Flvs.net? Your teacher isn't Ms. Huercamp is it?

OpenStudy (anonymous):

Yes and No

OpenStudy (anonymous):

flvsft.com

OpenStudy (anonymous):

I have Goeb

OpenStudy (anonymous):

Omg I remember her!! You should tell her Liam Lohrenz said hi :p I was a pain in the a$$ :DD but you should ask her in a live lesson cause when it comes time for the DBA you're going to have some trouble

OpenStudy (anonymous):

I know its my first time asking lol I do my own work

OpenStudy (anonymous):

I just am stuck on three of the multiple choices I dont know what im doing wrong

OpenStudy (anonymous):

Which three and what did you get on them?

OpenStudy (anonymous):

4. [9.01] What is the vertex for the graph of y = −2x2 + 8x − 13? (4 points) (2, 11) (2, −5) (−2, −37) (−2, −21) I know you use x=-b/2a then plug your answer into the x's but I dont get that when I do the problem

OpenStudy (anonymous):

5. [9.01] The graph of which equation is shown below? https://www.connexus.com/content/media/527934-762011-84933-AM-1383066450.png (4 points) y = −x2 + 6x − 5 y = −x2 − 6x − 5 y = x2 + 6x − 5 y = x2 − 6x − 5

OpenStudy (anonymous):

I haven't seen this stuff in forever o.o I'm in Geometry B right now... and this stuff is all gibberish.

OpenStudy (anonymous):

lol ya I know the formulas but I keep getting an answer thats not there :/ and I want to keep up my grade soo